An insect trap catches an average of four insects per 20 minute. Assuming that insects are caught randomly and independently, what is the probability that in a randomly chosen 20 minute period:
A) exactly five insects will be caught?
B) no more than two insects will be caught?
c) at least four insects will be caught?
D) over a randomly chosen one hour period what is the probability that exactly 10 insects will be caught?
The body temperature of a population of mammals are known to be normally distributed with a mean of 37.2 degrees and a standard deviation of 0.35 degress.
A) what proportion of the population have a tempreture of
i) more than 39 degrees ?
ii) less than 36.8 degrees?
iii) between 36.5 degrees and 38.5 degrees?
I will be very grateful if you could help me with these questions please.
Thank you very much
Birgul....

Answer by khwang(438)

Looks like the 1st question is related to Poission Distribution.
Every 20 min, lambda = 4. Let the random varible be X with
(I) Poission Distribution with mean Lambda = 4.
Pr(X =k) = (Lambda^k/k!) *e^(-Lambda)
A) exactly five insects will be caught?
B) no more than two insects will be caught?
c) at least four insects will be caught?
D) over a randomly chosen one hour period what is the probability that exactly
10 insects will be caught?
Sol: A) Lambda = 4, k= 5, Pr(X = 5) = (4^5/5!) *e^(-4) = 0.156
B) Lambda = 4, Pr(X <= 2) = 0.238
C) Lambda = 4, Pr(X => 4) = 1 - Pr(X <= 3) = 0.567
D) When time = 1 hr = 3*20 min, the mean Lambda is 3*4 = 12 .
Lambda = 12, Pr(X = 10) = 0.105
(II) Mu mean = 37.2 and a standard deviation sigma = 0.35 degress.
Convert it to the dstandard normal distribution by
z = (X- 37.2)/0.35 and look up the table to find the probability
i) more than 39 degrees ?
X = 39, Pr(X> 39)=1- NORMDIST(39,37.2,0.35,1) = 1.35541E-07

ii) less than 36.8 degrees?
X = 36.8, Pr(X< 36.8)= NORMDIST(36.8,37.2,0.35,1) = 0.127
iii) between 36.5 degrees and 38.5 degrees?
Pr(36.5 < X< 38.5)= Pr( X< 38.5) - Pr( X< 36.5)
= 0.98
Suggestion : Use MS Excel or Tables in the textbook.
Kenny